OpenStudy (anonymous):

Solve. 2x + 6 ≥ –4 A. x ≤ –8 B. x ≥ –5 C. x ≤ –5 D. x ≥ –8

OpenStudy (the_fizicx99):

subtract 6 from both sides. 2x > -10 /2 /2 x > -5 B

OpenStudy (anonymous):

ok so first subtract 6 from both sides so you get: \[2x \ge -10\] now we are going to divide by -10. When you divide an inequality by a negative number remember to flip the inequality symbol so if we divide by -10 we get: \[x \ge 5\]

OpenStudy (anonymous):

i mean \[x \ge -5\]

OpenStudy (anonymous):

Thank you :D